A pipe is draining water from a tank. After 5 minutes there are 300 gallons of water in the tank. After 10 minutes there are 260 gallons of water in the tank. How long will it take for the tank to be empty?
Found 2 solutions by TimothyLamb, lwsshak3:
Answer by TimothyLamb(4379)

time differential between the first and second measurement:
dt = 10 - 5 = 5 minutes
---
volume differential between the first and second measurement:
dv = 260 - 300 = -40 gallons
---
drain rate:
r = dv/dt
r = -(-40)/5 = 8 gal/min
---
at the first measurement the drain had been open 5 minutes ...
how much water was drained in that time?
r = v/t
v = rt
v = 8*5
v = 40
---
the tank's initial water volume was: 300 + 40 = 340
---
time to empty the tank:
r = v/t
t = v/r
t = 340/8
t = 42.5 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by lwsshak3(11628)

A pipe is draining water from a tank. After 5 minutes there are 300 gallons of water in the tank. After 10 minutes there are 260 gallons of water in the tank. How long will it take for the tank to be empty?
***
After 5 min,300-260=40 gallons of water was drained from the tank.
Therefore, with 260 gallons in the tank, at a rate of 40 gal/5min,
the tank will empty after (260/40)*5=32.5 minutes